A 45.0 mA current is carried by a uniformly wound air-core solenoid with 435 turns, a 19.5 mm diameter, and 14.0 cm length.

(a) Compute the magnetic field inside the solenoid.

(b) Compute the magnetic flux through each turn.

(c) Compute the inductance of the solenoid.

(d) Which of these quantities depends on the current? (Select all that apply.)

magnetic field inside the solenoid

magnetic flux through each turn

inductance of the solenoid
Answer #1

a)

Field inside the solenoid is given by

B = uo n i

B = 4 pi * 10^-7* 435 / 0.14* 0.045

B = 1.756*10^-4 T

b)

Flux through each turn

Flux = B A = 1.756*10^-4* 3.14* (19.5/ 2* 10^-3)^2

Flux = 5.242*10^-8 Wb

c)

Flux = L i

435* 5.242*10^-8 = L * 45*10^-3

L = 5.067*10^-4 H

d)

Flux and field both depends on current through the solenoid.

Option (a) & (b) are correct

===========

Comment in case any doubt, will reply for sure . Goodluck